OpenStudy (anonymous):

Simplify: 11 1/2 x 11^3

OpenStudy (anonymous):

convert 11 1/2 to an improper fraction first;you get: 23/2...now multiply: 23/2 x 11^3 = 23/2 x 1331 = 15306.5 ans